Electric potential at any point is given by –4x+ 5y+ 15z V , then magnitude of the electric field is

Grade:12th pass

1 Answers

Vikas TU
6 years ago
 
Dear Student,
we know that E=-dv/dx
So we calculate in each direction electric field separately,
E=-(dv/dx,dv/dy,dv/dz)
Ex=-dV/dx
=>Ex=-d(–4x + 5y + 15z)/dx
=>Ex=4
Ey=-dV/dy
=>Ey=-d(–4x + 5y + 15z)/dy
=>Ey=-5
Ez=-dV/dz
=>Ez=-d(–4x + 5y + 15z)/dz
=>Ez=-15
E=4i-5j-15k
|E|=√(16+25+225) = 16.31 V/m
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)